# Language Models Use Lookbacks to Track Beliefs

### v2

@@ 1 Introduction
Theory of Mind (ToM), the ability to infer others’ mental states, is an essential aspect of social and collective intelligence (Premack & Woodruff, 1978; Riedl et al., 2021). Recent studies have established that LMs can solve some tasks requiring ToM reasoning (Street et al., 2024; Strachan et al., 2024a; Kosinski, 2024), while others have highlighted shortcomings (Ullman, 2023; Sclar et al., 2025; Shapira et al., 2024, inter alia). Previous studies primarily rely on behavioral evaluations, which do not shed light on the internal mechanisms by which LMs encode and manipulate representations of mental states to solve (or fail to solve) such tasks (Hu et al., 2025; Gweon et al., 2023).
In this work, we examine how LMs internally represent and track beliefs of characters, a core aspect of ToM (Dennett, 1981; Wimmer & Perner, 1983). A classic example is the Sally-Anne test (Baron-Cohen et al., 1985), which evaluates ToM in humans by assessing whether individuals can track conflicting beliefs: Sally’s belief, which diverges from reality because of missing information, and Anne’s belief, which is updated based on new observations. Our goal is to determine whether LMs learn a systematic solution to such tasks or rely on superficial statistical association.
We construct CausalToM, a dataset of simple stories involving two characters, each interacting with an object to change its state, with the possibility of observing one another. We then analyze the internal mechanisms that enable Llama-3-70B-Instruct and Llama-3.1-405B-Instruct (Grattafiori et al., 2024) to reason about and answer questions regarding the characters’ beliefs about the state of each object (for a sample story, see Section 3 and for the full prompt refer to Appendix A).
@@ 2 The Lookback Mechanism
That is, the LM can use attention to dereference the pointer and retrieve the payload present in the residual stream of the recalled token (which might contain aggregated information from previous tokens), bringing it to the residual stream of the lookback token. Specifically, the pointer at the lookback token forms an attention query vector, while the address at the recalled token forms a key vector. The pointer and address are not necessarily exact copies of the source reference, but they do have a high dot product after being transformed by a query or key attention matrix, respectively. Hence, a QK-circuit (Elhage et al., 2021) is established, forming a bridge from the lookback token to the recalled token. The LM uses this bridge to move the payload that contains information needed to complete the subtask through the OV-circuit.
@@ 3 Experimental Setup: Dataset, Models, and Methods
Dataset    Existing datasets for evaluating ToM capabilities of LMs are designed for behavioral testing and lack counterfactual pairs needed for causal analysis (Kim & Sundar, 2012). To address this problem, we construct CausalToM, a structured dataset of simple stories, where each story involves two characters, each interacting with a distinct object causing the object to take a unique state. For example: “Character1 and Character2 are working in a busy restaurant. To complete an order, Character1 grabs an opaque Object1 and fills it with State1. Then Character2 grabs another opaque Object2 and fills it with State2.” We then ask the LM to reason about one of the characters’ beliefs regarding the state of an object: “What does Character1 believe Object2 contains?” We analyze the LM’s ability to track characters’ beliefs in two distinct settings. (1) No Visibility, where both characters are unaware of each other’s actions, and (2) Explicit Visibility, where explicit information about whether a character can/cannot observe the other’s actions is provided, e.g., “Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.” We also provide general task instructions (e.g., answer unknown when a character is unaware); refer to Appendices A & B for the full prompt and additional dataset details. All subsequent experiments are conducted on 80 samples that the model answers correctly. We also demonstrate generalization of the mechanism to BigToM dataset (Gandhi et al., 2024) in Appendix K.
Models    Our experiments analyze Llama-3-70B-Instruct and Llama-3.1-405B-Instruct models in FP16 and INT8 precision, respectively, using NNsight (Fiotto-Kaufman et al., 2025). Results for Llama-3.1-405B-Instruct can be found in Appendix L. Both models demonstrate strong behavioral performance in the no-visibility and explicit-visibility settings. We do not examine smaller models, as they are unable to coherently solve the CausalToM task.
@@
Causal Mediation Analysis    Our goal is to develop a mechanistic understanding of how LMs reason about characters’ beliefs and answer related questions (Saphra & Wiegreffe, 2024). A key method for conducting causal analysis is interchange interventions (Vig et al., 2020; Geiger et al., 2020; Finlayson et al., 2021), in which the LM is run on paired examples: an original input $\mathbf{o}$ and a counterfactual input $\mathbf{c}$, and certain internal activations in the LM run on the original input are replaced with those computed from the counterfactual, a process also known as activation patching. We begin our analysis by tracing information flow from key input tokens to the final output, by performing interchange interventions on the residual vectors. Specifically, we construct an intervention dataset where $\mathbf{o}$ contains a question about the belief of a character not mentioned in the story, while the story in $\mathbf{c}$ includes the same queried character, as shown in Fig. 2. The expected outcome of this intervention is a change in the final output of $\mathbf{o}$ from unknown to a state token, such as beer. We conduct similar interchange interventions for object and state tokens (refer to Appendix C for details).
@@
Desiderata Based Patching via Causal Abstraction    The causal mediation experiments provide a coarse-grained analysis of where information flows, but do not identify what information is being transferred. In a transformer, the first layer represents the input and the last layer represents the output, but we wish to know: what is represented in the middle? We analyze the internal mechanism using Causal Abstraction (Geiger et al., 2021; 2024); First, we hypothesize a high-level causal model of the computational steps from input to output (Sec. 4), and then align its variables with the LM’s internal activations (Sec. 5). We test the alignment through targeted interchange interventions on causal variables in the hypothesized model and hidden activations in the LM. If the LM produces the same output as the causal model under these aligned interventions, it provides evidence supporting the hypothesized causal model. We quantify this effect using interchange intervention accuracy (IIA; Geiger et al., 2022), which measures the proportion of cases where the intervened causal model and intervened LM agree. See Appendix D for more details.
In addition to measuring IIA on entire residual stream vectors, we also intervene on localized subspaces to further isolate causal variables. To identify the subspace of a specific variable, we employ Desiderata-based Component Masking (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024). This method learns a sparse binary mask over the activation space that maximizes the logit of the hypothesized causal model output. We train a mask to select singular vectors of the activation space that encode a high-level variable (see Appendix F for details). Our experiments in Sec. 5 report both interventions on the full residual stream and on the identified subspaces.
@@ 4 Hypothesized High-Level Causal Model of Belief Tracking
Here we start with an overview of our hypothesized causal model of belief-tracking when characters are not aware of each other’s actions. The causal model is an algorithmic process that has variables with structural roles that do not refer to the details of a transformer architecture. Appendix E presents the full pseudocode of the causal model. In Section 5, we will present experiments to verify that the causal model’s variables align with representations in the transformer.
Belief tracking begins when the causal model assigns ordering IDs (OIs; [Uncaptioned image], [Uncaptioned image], [Uncaptioned image]) to each character, object, and state token, marking their order of appearance. For instance, in the example in Fig. 3, Bob is assigned first character OI ([Uncaptioned image]), and Carla is assigned the second character OI([Uncaptioned image]). Then it uses these OIs in two lookback mechanisms:
@@ 5 Verifying the Hypothesized Causal Model of Belief Tracking
Because the last step of the causal model is easiest to understand, we proceed through the experiments in reverse order, beginning with an experiment to verify the final “answer lookback” stage. After this instructive starting point, we work backward to verify the earlier steps of the model. Additional results can be found in Appendix G and H.
@@ 5.2 Step i: Binding Lookback – Linking Characters, Objects, and States
In the LM, we interchange the residual streams of the character and object tokens layer-by-layer, while keeping the residual stream of the state token fixed. As shown in Fig. 6b, this experiment reveals alignment between layers $20$ and $34$, indicating that source reference is encoded in the residual streams of the character and object tokens within this layer range. Additional results are provided in Appendix G, where Fig. 13 shows that freezing the residual stream of the state token is necessary for this alignment to emerge. These findings support our hypothesis that source reference is present in the character and object tokens and is subsequently transferred to the recalled and lookback tokens.
@@
Localizing the Pointer Information    Finally, we localize the pointer copies of the character and object OIs to their corresponding tokens in the question and to the final token. See Appendices G &  H for details of the experiments and results.
@@ 6 Impact of Visibility Conditions on Belief Tracking Mechanism
Figure 7: Visibility Lookback When one (observing) character can see another (observed) character, the LM assigns a visibility ID (Refer to caption) to the visibility sentence where this relation is defined. An address copy of this visibility ID remains in the visibility sentence’s residual stream. A pointer copy of the visibility ID is transferred to the subsequent tokens’ residual stream. The LM dereferences this pointer through a QK-circuit, bringing forward the payload ( Refer to caption ), when processing subsequent tokens. Based on initial evidence, this payload contains the observed character’s OI(Refer to caption). See Appendix I for details. This mechanism allows the model to incorporate the observed character’s knowledge into the observing character’s belief state, enabling more complex belief reasoning.
Although our two-character setting is unable to discern the exact semantics of the payload in the visibility lookback, our observations are consistent with the payload encoding the observed character’s OI. Our initial observations suggest another lookback where the story sentence associated with the observed character serves as the source reference, and its payload encodes information about the observed character. The observed character’s OI appears to be retrieved by the lookback tokens of the Visibility lookback, with causal effects on the queried character’s awareness (see App. I for details).
@@ 7 Related Work
Theory of mind in LMs    Theory of mind in LMs has been widely benchmarked (Le et al., 2019; Shapira et al., 2023; Wu et al., 2023; Kim et al., 2023; Xu et al., 2024; Jin et al., 2024; Chan et al., 2024; Strachan et al., 2024b). However, these benchmarks lack adequate counterfactuals for the binding manipulations we need, so we made CausalToM (Section B).
Entity tracking in LMs    Entity tracking and variable binding are crucial abilities for LMs to exhibit not only coherent ToM capabilities, but also neurosymbolic reasoning. Many existing works have attempted to decipher this ability in LMs (Li et al., 2021; Davies et al., 2023; Feng & Steinhardt, 2023; Kim & Schuster, 2023; Prakash et al., 2024; Feng et al., 2024; Dai et al., 2024; Wu et al., 2025). Our work builds on their empirical insights and extends the current understanding of how LMs bind various entities defined in context.
Mechanistic interpretability of theory of mind    Few studies explored the underlying mechanisms of ToM of LM (Zhu et al., 2024; Bortoletto et al., 2024; Herrmann & Levinstein, 2024). Those studies use probing (Alain, 2016; Belinkov, 2022) to identify internal representations of beliefs and steering (Rimsky et al., 2023; Li et al., 2024) to control LMs by manipulating their activations. However, the mechanism by which LMs solve those tasks remains a black box, limiting our ability to understand, predict, and control LMs’ behaviors.
@@ 9 Ethics Statement
This work involves experiments conducted exclusively on synthetic text generated by the authors. No human subjects, personal data, or sensitive user information were collected or analyzed. The models we analyze are publicly released LLMs (Llama-3-70B-Instruct and Llama-3.1-405B-Instruct). All interventions and analyses were performed locally without querying proprietary APIs. We acknowledge that research on ToM in LMs may be misinterpreted as suggesting human-like cognition or intentionality. We explicitly caution that our findings demonstrate internal computational mechanisms, not conscious reasoning, and should not be construed as evidence of sentience or moral agency in LMs. Our causal intervention techniques reveal latent structures within pretrained models but do not modify the underlying weights or enable novel capabilities. Nevertheless, reverse-engineering latent belief representations could, in principle, be misused for behavioral steering or manipulation. To mitigate this, we provide our code strictly for transparency and further scientific audit, rather than for application in deployed systems.
@@ 10 Reproducibility Statement
To facilitate reproducibility, we release the full CausalToM dataset, including all story templates and the code used to generate the various story instances that serve as counterfactual variants in our experiments. The repository, that can be accessed at https://belief.baulab.info, contains all scripts required to construct the dataset, extract activations, perform interchange interventions, and compute causal mediation metrics, along with the hyperparameters and random seeds used for subspace identification via DCM. All experiments were conducted using publicly available open-weight models (Llama-3-70B-Instruct and Llama-3.1-405B-Instruct). The interchange intervention experiments can be reproduced either by hosting these models locally or by using remotely hosted instances via NDIF (Fiotto-Kaufman et al., 2025), if sufficient local compute is not available.

### v3

@@ 1 Introduction
Theory of Mind (ToM), the ability to infer others’ mental states, is an essential aspect of social and collective intelligence (Premack and Woodruff, 1978; Riedl et al., 2021). Recent studies have established that LMs can solve some tasks requiring ToM reasoning (Street et al., 2024; Strachan et al., 2024a; Kosinski, 2024), while others have highlighted shortcomings (Ullman, 2023; Sclar et al., 2025; Shapira et al., 2024, inter alia). Previous studies primarily rely on behavioral evaluations, which do not shed light on the internal mechanisms by which LMs encode and manipulate representations of mental states to solve (or fail to solve) such tasks (Hu et al., 2025; Gweon et al., 2023).
In this work, we examine how LMs internally represent and track beliefs of characters, a core aspect of ToM (Dennett, 1981; Wimmer and Perner, 1983). A classic example is the Sally-Anne test (Baron-Cohen et al., 1985), which evaluates ToM in humans by assessing whether individuals can track conflicting beliefs: Sally’s belief, which diverges from reality because of missing information, and Anne’s belief, which is updated based on new observations. Our goal is to determine whether LMs learn a systematic solution to such tasks or rely on superficial statistical association.
We construct CausalToM, a dataset of simple stories involving two characters, each interacting with an object to change its state, with the possibility of observing one another. We then analyze the internal mechanisms that enable Llama-3-70B-Instruct, Llama-3.1-405B-Instruct and Qwen2.5-14B-Instruct (Grattafiori et al., 2024; Bai et al., 2023) to reason about and answer questions regarding the characters’ beliefs about the state of each object (for a sample story, see Section 3 and for the full prompt refer to Appendix A).
@@ 2 The Lookback Mechanism
That is, the LM can use attention to dereference the pointer and retrieve the payload present in the residual stream of the recalled token (which might contain aggregated information from previous tokens), bringing it to the residual stream of the lookback token. Specifically, the pointer at the lookback token forms an attention query vector, while the address at the recalled token forms a key vector. The pointer and address are not necessarily exact copies of the source reference, but they do have a high dot product after being transformed by a query or key attention matrix, respectively. Hence, a QK-circuit (Elhage et al., 2021) is established, forming a bridge from the lookback token to the recalled token. The LM uses this bridge to move the payload that contains information needed to complete the subtask through the OV-circuit. See Appendix C for background on the residual-stream framework as well as on QK- and OV-circuits.
@@ 3 Experimental Setup: Dataset, Models, and Methods
3.1 Dataset
Existing datasets for evaluating ToM capabilities of LMs are designed for behavioral testing and lack counterfactual pairs needed for causal analysis (Kim and Sundar, 2012). To address this problem, we construct CausalToM, a structured dataset of simple stories, where each story involves two characters, each interacting with a distinct object causing the object to take a unique state. For example: “Character1 and Character2 are working in a busy restaurant. To complete an order, Character1 grabs an opaque Object1 and fills it with State1. Then Character2 grabs another opaque Object2 and fills it with State2.” We then ask the LM to reason about one of the characters’ beliefs regarding the state of an object: “What does Character1 believe Object2 contains?” We analyze the LM’s ability to track characters’ beliefs in two distinct settings. (1) No Visibility, where both characters are unaware of each other’s actions, and (2) Explicit Visibility, where explicit information about whether a character can/cannot observe the other’s actions is provided, e.g., “Bob can observe Carla’s actions. Carla cannot observe Bob’s actions.” We also provide general task instructions (e.g., answer unknown when a character is unaware); refer to Appendices A & B for the full prompt and additional dataset details. All subsequent experiments are conducted on 80 samples that the model answers correctly. We also demonstrate generalization of the mechanism to BigToM dataset (Gandhi et al., 2024) in Appendix M.
3.2 Models
Our experiments analyze Qwen2.5-14B-Instruct, Llama-3-70B-Instruct, and Llama-3.1-405B-Instruct models in FP32, FP16, and INT8 precision, respectively, using NNsight (Fiotto-Kaufman et al., 2025). Results for Qwen2.5-14B-Instruct and Llama-3.1-405B-Instruct can be found in Appendix N. We selected these models due to their strong behavioral performance under both no-visibility and explicit-visibility conditions. Additional details can be found in Appendix D.
3.3 Causal Mediation Analysis
Our goal is to develop a mechanistic understanding of how LMs reason about characters’ beliefs and answer related questions (Saphra and Wiegreffe, 2024). A key method for conducting causal analysis is interchange interventions (Vig et al., 2020; Geiger et al., 2020; Finlayson et al., 2021), in which the LM is run on paired examples: an original input $\mathbf{o}$ and a counterfactual input $\mathbf{c}$, and certain internal activations in the LM run on the original input are replaced with those computed from the counterfactual, a process also known as activation patching. We begin our analysis by tracing information flow from key input tokens to the final output, by performing interchange interventions on the residual vectors. Specifically, we construct an intervention dataset where $\mathbf{o}$ contains a question about the belief of a character not mentioned in the story, while the story in $\mathbf{c}$ includes the same queried character, as shown in Fig. 2. The expected outcome of this intervention is a change in the final output of $\mathbf{o}$ from unknown to a state token, such as beer. We conduct similar interchange interventions for object and state tokens (refer to Appendix E for details).
@@ 3.3 Causal Mediation Analysis
Desiderata Based Patching via Causal Abstraction    The causal mediation experiments provide a coarse-grained analysis of where information flows, but do not identify what information is being transferred. In a transformer, the first layer represents the input and the last layer represents the output, but we wish to know: what is represented in the middle? We analyze the internal mechanism using Causal Abstraction (Geiger et al., 2021; 2024); First, we hypothesize a high-level causal model of the computational steps from input to output (Sec. 4), and then align its variables with the LM’s internal activations (Sec. 5). We test the alignment through targeted interchange interventions on causal variables in the hypothesized model and hidden activations in the LM. If the LM produces the same output as the causal model under these aligned interventions, it provides evidence supporting the hypothesized causal model. We quantify this effect using interchange intervention accuracy (IIA; Geiger et al., 2022), which measures the proportion of cases where the intervened causal model and intervened LM agree. See Appendix F for more details.
In addition to measuring IIA on entire residual stream vectors, we also intervene on localized subspaces to further isolate causal variables. To identify the subspace of a specific variable, we employ Desiderata-based Component Masking (De Cao et al., 2020; Davies et al., 2023; Prakash et al., 2024). This method learns a sparse binary mask over the activation space that maximizes the logit of the hypothesized causal model output. We train a mask to select singular vectors of the activation space that encode a high-level variable (see Appendix H for details). Our experiments in Sec. 5 report both interventions on the full residual stream and on the identified subspaces.
@@ 4 Hypothesized High-Level Causal Model of Belief Tracking
Here we start with an overview of our hypothesized causal model of belief-tracking when characters are not aware of each other’s actions. The causal model is an algorithmic process that has variables with structural roles that do not refer to the details of a transformer architecture. Appendix G presents the full pseudocode of the causal model. In Section 5, we will present experiments to verify that the causal model’s variables align with representations in the transformer.
Belief tracking begins when the causal model assigns ordering IDs (OIs; [Uncaptioned image], [Uncaptioned image], [Uncaptioned image]) to each character, object, and state token, marking their order of appearance. For instance, in the example in Fig. 3, Bob is assigned first character OI ([Uncaptioned image]), and Carla is assigned the second character OI([Uncaptioned image]). Please refer to Appendix C for more information about OI. Then it uses these OIs in two lookback mechanisms:
@@ 5 Verifying the Hypothesized Causal Model of Belief Tracking
Because the last step of the causal model is easiest to understand, we proceed through the experiments in reverse order, beginning with an experiment to verify the final “answer lookback” stage. After this instructive starting point, we work backward to verify the earlier steps of the model. Additional results can be found in Appendix I and J.
@@ 5.2 Step i: Binding Lookback – Linking Characters, Objects, and States
In the LM, we interchange the residual streams of the character and object tokens layer-by-layer, while keeping the residual stream of the state token fixed. As shown in Fig. 6b, this experiment reveals alignment between layers $20$ and $34$, indicating that source reference is encoded in the residual streams of the character and object tokens within this layer range. Additional results are provided in Appendix I, where Fig. 13 shows that freezing the residual stream of the state token is necessary for this alignment to emerge. These findings support our hypothesis that source reference is present in the character and object tokens and is subsequently transferred to the recalled and lookback tokens.
Localizing the Pointer Information    Finally, we localize the pointer copies of the character and object OIs to their corresponding tokens in the question and to the final token. See Appendices I &  J for details of the experiments and results.
@@ 6 Impact of Visibility Conditions on Belief Tracking Mechanism
Figure 7: Visibility Lookback When one (observing) character can see another (observed) character, the LM assigns a visibility ID (Refer to caption) to the visibility sentence where this relation is defined. An address copy of this visibility ID remains in the visibility sentence’s residual stream. A pointer copy of the visibility ID is transferred to the subsequent tokens’ residual stream. The LM dereferences this pointer through a QK-circuit, bringing forward the payload ( Refer to caption ), when processing subsequent tokens. Based on initial evidence, this payload contains the observed character’s OI(Refer to caption). See Appendix K for details. This mechanism allows the model to incorporate the observed character’s knowledge into the observing character’s belief state, enabling more complex belief reasoning.
@@
Although our two-character setting is unable to discern the exact semantics of the payload in the visibility lookback, our observations are consistent with the payload encoding the observed character’s OI. Our initial observations suggest another lookback where the story sentence associated with the observed character serves as the source reference, and its payload encodes information about the observed character. The observed character’s OI appears to be retrieved by the lookback tokens of the Visibility lookback, with causal effects on the queried character’s awareness (see App. K for details).
@@ 7 Related Work
Theory of mind in LMs    Theory of mind in LMs has been widely benchmarked (Le et al., 2019; Shapira et al., 2023; Wu et al., 2023; Kim et al., 2023; Xu et al., 2024; Jin et al., 2024; Chan et al., 2024; Strachan et al., 2024b). However, these benchmarks lack adequate counterfactuals for the binding manipulations we need, so we made CausalToM (Section B). Although several methods have been proposed to improve ToM performance in LMs (e.g., Sclar et al., 2023; Moghaddam and Honey, 2023; Zhou et al., 2023; Wilf et al., 2024; Hou et al., 2024), our goal is not to enhance ToM abilities. Instead, we aim to understand the existing ToM behaviors of LMs by reverse-engineering the underlying mechanisms that support belief tracking.
Entity tracking in LMs    Entity tracking and variable binding are essential for LMs to support not only coherent ToM reasoning but also broader neurosymbolic inference. A substantial body of work has sought to uncover how LMs implement these abilities. Li et al. (2021) demonstrated that LMs form internal representations that encode the dynamic states of entities, which can be identified and manipulated using linear probes (Belinkov, 2022). Building on this, Davies et al. (2023), Prakash et al. (2024) and Yang et al. (2025) found that modern LMs rely on a small set of attention heads that track an entity’s positional information to retrieve its corresponding attributes. Complementing these findings, Feng and Steinhardt (2023) showed that LMs encode similar relational information in the form of Binding IDs within their hidden representations. Extending this line of work, Dai et al. (2024) demonstrated that LMs also learn more symbolic representation, referred to as Ordering IDs, that capture the positional information in low rank subspace of a token’s residual stream. Wu et al. (2025) trained a Transformer on symbolic programs and showed that variable binding emerges over three distinct training phases, progressing from shallow heuristics to a fully systematic mechanism. Our work builds on this growing literature by uncovering the end-to-end mechanism that enables variable binding and by further elucidating how ToM-related reasoning is implemented through these binding processes within LM internal representations.
Mechanistic interpretability of theory of mind    Recently, a few studies have begun probing the internal representations underlying ToM abilities in LMs. Zhu et al. (2024) showed that the belief states of different agents can be linearly decoded using simple probes, and that intervening on these representations leads to substantial changes in ToM performance. Building on this, Herrmann and Levinstein (2024) examined how prompt variation and fine-tuning affect the stability and structure of belief representations. Additionally, Bortoletto et al. (2024) proposed adequacy criteria, such as accuracy, coherence, uniformity, and use, for determining when an internal representation in an LLM should count as belief-like. While these works provide valuable insight into how belief information is encoded, they do not illuminate the mechanisms by which LMs actually solve ToM tasks, limiting our ability to understand, predict, and control model behavior.
@@ 9 Ethics Statement
We recognize that ToM reasoning can be misused for deception, persuasion, or targeted advertising. For this reason, our work deliberately avoids amplifying such capabilities and instead focuses solely on measuring and characterizing the ToM abilities that models already possess. Our goal is not to accelerate a model’s capacity for ToM reasoning, but to understand the underlying mechanisms that would allow us to detect ToM-related behavior by directly examining its neural fingerprint.
We believe that this kind of internal neural analysis is essential, particularly when there are concerns that models might learn to deceive or persuade. Because deception and persuasion are, by definition, difficult or impossible to identify from a model’s outputs alone, it is crucial to understand their neural signatures. Doing so enables us to distinguish between surface-level outputs that merely appear to reflect certain behaviors and internal reasoning pathways that genuinely reveal ToM processes.
Finally, we acknowledge that research on ToM in LMs can be misinterpreted as implying human-like cognition or intentionality. We explicitly caution that our findings describe internal computational mechanisms, not conscious reasoning, and should not be taken as evidence of sentience or moral agency in LMs.
@@ 10 Reproducibility Statement
To facilitate reproducibility, we release the full CausalToM dataset, including all story templates and the code used to generate the various story instances that serve as counterfactual variants in our experiments. The repository, which can be accessed at https://belief.baulab.info, contains all scripts required to construct the dataset, extract activations, perform interchange interventions, and compute causal mediation metrics, along with the hyperparameters and random seeds used for subspace identification via DCM. All experiments were conducted using publicly available open-weight models (Llama-3-70B-Instruct, Llama-3.1-405B-Instruct, and Qwen2.5-14B-Instruct) hosted on Huggingface. Experiments with Llama-3-70B-Instruct and Qwen2.5-14B-Instruct were carried out locally on two 80GB NVIDIA A100 GPUs, while experiments with Llama-3.1-405B-Instruct were conducted remotely using the NDIF (Fiotto-Kaufman et al., 2025) service. Results for Llama-3-70B-Instruct and Qwen2.5-14B-Instruct can also be reproduced using NDIF if sufficient local computing resources are unavailable.